OpenStudy (anonymous):

On a planet whose size is same and mass 4 times as that of our earth , find the amount of work done to lift 3 kg mass vertically upwards through 3 m distance on the planet. The value of g on the surface of earth is g=10 m/s^2

OpenStudy (mayankdevnani):

mass=4times of earth=4*3=12kg

OpenStudy (mayankdevnani):

ok @ashna

OpenStudy (mayankdevnani):

work done=potential energy

OpenStudy (mayankdevnani):

potential energy=mass*gravity*height

OpenStudy (mayankdevnani):

P.E=12*10*3=120*3=360J

OpenStudy (mayankdevnani):

so work done=360J

OpenStudy (mayankdevnani):

got it @ashna

OpenStudy (anonymous):

yea :) ty !

OpenStudy (mayankdevnani):

right @ajprincess